OpenStudy (anonymous):

what angles are coterminal with 160 degrees

OpenStudy (anonymous):

@mertsj can u help me

OpenStudy (anonymous):

anyone....?

jimthompson5910 (jim_thompson5910):

Anything coterminal to 160 degrees will be separated by 360 degrees. In other words, simply add/subtract 360 to/from 160 to get a new coterminal angle So one coterminal angle is 160 + 360 = 520 degrees

OpenStudy (anonymous):

so would one be 20?

jimthompson5910 (jim_thompson5910):

another coterminal angle would be 160 - 360 = -200 degrees

OpenStudy (anonymous):

thts not one of the options given to me. the options are : 20, 520 and -160, 200, and -200,-560, and 520, 700

OpenStudy (anonymous):

Coterminal angles Ac to angle A may be obtained by adding or subtracting k*360 degrees or k* (2 Pi). Hence Ac = A + k*360o if A is given in degrees. or Ac = A + k*(2 PI) if A is given in radians. where k is any negative or positive integer.

jimthompson5910 (jim_thompson5910):

sorry was delayed a bit Notice how -200 is one of our coterminal angles So if we subtract another 360 from that, we get -200 - 360 = -560 So the answer is choice C

OpenStudy (anonymous):

Find a positive and a negative angle coterminal with a 160° angle. 160° – 360° = –200° 160° + 360° = 520° A –200° angle and a 520° angle are co terminal with a 160° angle.

OpenStudy (anonymous):

can u help me out with a couple more pleaseee?

OpenStudy (anonymous):

its a different question, but its on trig

OpenStudy (anonymous):